OpenStudy (anonymous):

The regular price of a child's entry ticket to a water park is $6 less than that for an adult's. The park offers half off of all entry tickets during the off-peak season. The Jordans paid a total of $69 for 2 adult tickets and 1 child's ticket to the water park during the off-peak season. The following equation represents this situation, where x represents the regular price of an adult ticket. 69 = x + 1/2 (x - 6) What is the regular price of a child's ticket?

OpenStudy (anonymous):

69 = x + 1/2 (x - 6) solve for x: 69 = x +.5X -3 72 = 1.5x x = 72/1.5 = 48 = regular price of adult ticket (this is given in the question) regular price of child ticket = adult ticket - 6 = 48 - 6 = $42